Fermi National Laboratory


Accelerator Update for June 22, 2005 - June 24, 2005

Wednesday June 22
The day shift began with Operations monitoring store 4221 and the stash, with Tevatron (TeV) experts conducting an end-of-store study, with the Antiproton Source (Pbar) stacking, and with the MiniBooNE, and NuMI, and Meson MCenter experiments taking beam. Operations began sending beam to the Meson MTest experiment at 8:49 AM.

The TeV end-of-store study ended at 9:40 AM.

Operations terminated store 4221 at 9:43 AM.

The D-Zero (D0) and Collider Detector at Fermilab (CDF) experimenters accessed their collision halls at 10:02 AM. Both experiments were out of access by 10:41 AM.

At 10:56 AM, MiniBooNE experimenters reported that they were having data acquisition problems and asked to have their beam halted.

MiniBooNE came back on at 12:22 PM.

Operations put the TeV through a dry squeeze at 12:23 PM.

Operations put the TeV into shot setup at 12:54 PM.

At 2:56 PM, shortly after the last antiprotons were loaded the TeV suffered a sector F3 quench.

Operators reestablished beam to MCenter at 3:42 PM.

At 4:09 PM, a Linac RF station (KRF4) started having a phase problem due to a LCW valve position. An expert investigated.

Operations put the TeV through a dry squeeze at 6:18 PM.

Pond Water At 8:08 PM, Operations reported that the I- Source had tripped off for the first of many times due to a high voltage problem. An expert came in and reported that it was running okay at 9:42 PM.

At 8:37 PM, Operations reported that the TeV suffered a sector B quench. TeV experts were conducting a low beta magnet study at the time.

At 9:13 PM, Operations reported that a TeV sector B1 pond water pump developed a leak. The duty mechanic responded and verified a leak.

Thursday June 23
The midnight shift began with Operations monitoring a Recycler stash, with Pbar stacking, with the TeV in a dry squeeze, and with MiniBooNE and NuMI taking beam.

Operations put the TeV into shot setup at 12:40 AM.

Operations established combination store 4228 at 2:57 AM with an initial luminosity of 83.95E30. Pbar resumed stacking.

MCenter resumed taking beam at 3:04 AM.

At 3:31 AM, Operations reported having problems with MECAR. An operator called an expert. The expert recommended rebooting the MECAR system, which Operations did. Operators also rebooted the MIBEAM. MI beam resumed at 4:20 AM.

Operations notified a water expert of the B1 pond pump leak at 7:32 AM. The expert started his repairs.

Operations put the Recycler into shot setup to stash antiprotons at 8:40 AM.

The stash ended and MCenter resumed taking beam at 10:14 AM.

Operations prepared the Pbar tunnels for an access at 10:16 AM. Experts found a ground fault on A:LQ.

MTest requested beam at 10:41 AM, but operators couldn’t find the setting for the 4 GeV beam they want. An operator called in an expert.

At 12:50 PM, Operations reported that store 4228 aborted due to a B0 quadrupole (BOQ3) tripping off. There was no quench. Experts discovered that the problem wasn’t with B0Q3 and began trouble shooting the QPM equipment.

A MI expert began a slipstacking study at 2 PM.

Around 2:20 PM, Operations reported that KRF3 had tripped off. After operators got the station back on, Booster couldn’t hold beam. Operations reported at 3:26 PM, that they found a phase problem with KRF7 that was effecting Booster injection.

TeV expert reported at 4:50 PM that they had finished their B0 and D0 QPM work. Operations put the TeV into a dry squeeze.

TeV experts began a study period at 6:31 PM.

At 8:28 PM, the NuMI horn tripped off due to a LCW flow indication. Operators called in water experts.

At 9:25 PM, the 400 MeV area chopper tripped off and wouldn’t reset. Operators investigated and then called in experts. The expert reported at 10:20 PM that the trip was due to a LCW flow fault. The chopper was back on by 10:21 PM.

At 10:39 PM, Operations reported a power glitch and the following trips: Booster RF, the MI power supply loop, all the cryo system wet engines and helium compressors, Linac RF, the Pbar Debuncher, the TeV anode, and numerous other systems. Operators called in experts.

At 10:55 PM, the Duty Electrician reported that feeder 46 had tripped.

The Booster expert completed his work on the 400 MeV chopper at 11:01 PM.

Friday June 24
The midnight shift began with Operations and machine experts working to recover from the feeder 46 failure. Operations monitored the stack and stash, since neither machine was affected by the glitch.

At 12:20 AM, the 400 MeV chopper tripped off again. Experts investigated.

The High Voltage expert reported at 12:47 AM that they had found the feeder 46B fault between TeV sectors D0 and F0.

Operations reported establishing beam to the Linac dump at 1:39 AM.

The High Voltage expert reported at 1:40 AM that the power to feeder 46B had been restored.

Operations reported at 2:16 AM that the NuMI horn was back up. The expert said it wasn’t a water problem, but a bad air filter on a vent blower.

Operations reestablished beam to MiniBooNE at 2:18 AM.

Operations established beam to MI at 2:23 AM.

At 2:47 AM, the Debuncher lithium lens power supply tripped off. Operators called in an expert.

Operations established beam to NuMI at 2:56 AM.

The lithium lens power supply work finished and Operations established beam to the Accumulator at 5:41 AM. Pbar resumed stacking.

Around 8 AM, Operations reported the following: TRF1 was tripping—operators notified an expert, and operators were trying to get beam to MTest, but couldn’t see the beam on the SWICs—operators contacted an expert.

The Plan for Friday
The plans for today are to put the TeV through a dry squeeze and then a wet squeeze, followed by shot setup. The plans for the weekend are to stack, stash, and store.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 6/24/2005   email Fermilab